# PTR: Precision-Driven Tool Recommendation for Large Language Models

### v1

@@ 5.1 Implementation Details
Baselines. We considered the following baselines: Random, which randomly select from historical tools; BM25 (Robertson et al., 2009), a classical sparse retrieval method that extends TF-IDF by leveraging term frequency and inverse document frequency of keywords; Contriever (Izacard et al., 2021), which utilizes inverse cloze tasks, cropping for positive pair generation, and momentum contrastive training to develop dense retrievers; SBERT (Reimers & Gurevych, 2019), a library providing BERT-based sentence embeddings. Specifically, we use all-mpnet-base-v2; TAS-B (Hofstätter et al., 2021), the retriever introduces an efficient topic-aware query and balanced margin sampling technique; And SimCSE (Gao et al., 2021), a simple contrastive learning framework that greatly advances state-of-the-art sentence embeddings.
@@ 5.3 Further Analysis
In this section, we conduct an in-depth analysis of the effectiveness for PTR, using the same datasets and evaluation metrics. The results are presented in Table 3.
Table 3: Ablation study of Tool Bundle Acquisition (w/o Tool Bundle Acquisition).
Methods (w/o Tool Bundle Acquisition) ToolLens MetaTool RecTools
Recall@K NDCG@K TRACC Recall@K NDCG@K TRACC Recall@K NDCG@K TRACC
PTR + open-mistral-7b 0.221 0.264 0.171 0.695 0.882 0.612 0.532 0.912 0.270
PTR + GPT-3.5-turbo 0.264 0.381 0.208 0.724 0.919 0.656 0.541 0.913 0.430
PTR + GPT-4o 0.283 0.391 0.235 0.745 0.922 0.677 0.581 0.916 0.439
w/o Tool Bundle Acquisition. This variant omits the tool bundle acquisition stage, resulting in queries being exclusively mapped to unresolved problems without any existing recommended tools. The observed decline in performance for this variant further supports the effectiveness of tool bundles in identifying potential recommended tools, thereby refining the unresolved problems and achieving precise tool recommendations. Moreover, as illustrated in Table 3, the random approach alone is largely ineffective for tool recommendations. However, as presented in Table 2, when combined with functional coverage mapping and multi-view-based re-ranking, the final recommendation performance improves significantly. This underscores the importance of the last two recommendation stages.
@@ 6 Related work
6.1 Recommendation for LLMs
Recent research has explored a variety of recommendation techniques to enhance Large Language Models (LLMs), integrating capabilities across multiple dimensions. Data recommendation (Xu et al., 2020; Ouyang et al., 2022) is crucial for selecting relevant datasets to fine-tune models for specific domains, ensuring ongoing performance improvements. Memory recommendation (Borgeaud et al., 2022; Gao & Zhang, 2024a) facilitates the retrieval of relevant past experiences or interactions, improving continuity, consistency, and long-term context in multi-turn conversations. Knowledge base recommendation (Gao et al., 2023; Hu et al., 2023; Petroni et al., 2019; Lewis et al., 2020) enhances factual grounding by retrieving the most pertinent information from external sources, ensuring that model outputs are accurate and up to date. Architecture recommendation (Elsken et al., 2019; Fedus et al., 2022) optimizes model performance by dynamically selecting the most appropriate model components or layers to activate for different tasks, thereby improving efficiency. Lastly, prompt recommendation (Shin et al., 2020; Reynolds & McDonell, 2021; Li & Liang, 2021; Wang et al., 2022; Liu et al., 2023) guides LLMs in utilizing the most effective input prompts, thereby enhancing the quality of generated responses through optimized input-output interactions. Together, these recommendation techniques form a comprehensive framework that enhances the adaptability, efficiency, and task-specific performance of LLMs. However, there remains a lack of research on tool recommendation. In this work, we motivate to seek to provide a clear definition of tool recommendation and proposes an effective recommendation method. Additionally, new datasets and metrics are created to advance research in this area.
6.2 Tool Retrieval
Initially, term-based methods such as BM25 (Robertson et al., 2009) and TF-IDF (Sparck Jones, 1972) were employed to measure the similarity between queries and tool documents by identifying exact term matches. Subsequently, with the development of dense retrievers (Karpukhin et al., 2020; Guu et al., 2020; Xiong et al., 2020), the semantic relationships between queries and tool descriptions have been more effectively captured through neural networks. Recently, novel approaches for training retrievers have emerged. For example, Confucius (Gao et al., 2024) selects tools by defining three levels of scenarios, ranging from easy to difficult, to train and deepen the LLM’s understanding of tools. Additionally, execution feedback is iteratively utilized to refine the tool selection process (Wang et al., 2023; Xu et al., 2024). Furthermore, ToolkenGPT (Hao et al., 2024) enhances tool selection by representing each tool as a token (”toolken”) and learning an embedding for it, thereby enabling tool calls in the same manner as generating regular word tokens. Moreover, some research has focused on addressing the diversity of retrieval (Carbonell & Goldstein, 1998; Gao & Zhang, 2024b), which can effectively enhance the quality of multiple tools used by query. Despite their comprehensive nature, tool retrieval systems present notable limitations. The inclusion of superfluous tools can introduce noise, thereby interfering with the LLM’s performance and task execution, and these systems are often unable to dynamically adjust the toolset. In this work, we extend our approach beyond getting a rough toolset by ensuring that the tools in the recommended toolset are as accurate as possible in terms of both quality and quantity.

### v2

@@ 4.1 Implementation Details
Baselines. We considered the following baselines: Random, which randomly select from historical tools; BM25 (robertson2009probabilistic), a classical sparse retrieval method that extends TF-IDF by leveraging term frequency and inverse document frequency of keywords; Contriever (izacard2021unsupervised), which utilizes inverse cloze tasks, cropping for positive pair generation, and momentum contrastive training to develop dense retrievers; SBERT (reimers-2019-sentence-bert), a library providing BERT-based sentence embeddings. Specifically, we use all-mpnet-base-v2; TAS-B (hofstatter2021efficiently), the retriever introduces an efficient topic-aware query and balanced margin sampling technique; And SimCSE (gao2021simcse), a simple contrastive learning framework that greatly advances state-of-the-art sentence embeddings.
@@ 4.3 Further Analysis
In this section, we conduct an in-depth analysis of the effectiveness for PTR, using the same datasets and evaluation metrics.
Dataset Variant Recall@K NDCG@K TRACC
ToolLens w/o Stage-1 0.283 0.391 0.235
w/o Stage-2 0.355 0.455 0.322
w/o Stage-3 0.351 0.448 0.318
PTR 0.368 0.467 0.336
MetaTool w/o Stage-1 0.745 0.922 0.677
w/o Stage-2 0.762 0.945 0.678
w/o Stage-3 0.759 0.942 0.676
PTR 0.774 0.956 0.690
RecTools w/o Stage-1 0.581 0.916 0.439
w/o Stage-2 0.598 0.888 0.552
w/o Stage-3 0.601 0.892 0.566
PTR 0.609 0.896 0.591
Table 2: Ablation results on three datasets.
@@
Table 2 reports the ablation results of PTR, where we individually remove each of the three stages, Tool Bundle Acquisition, Functional Coverage Mapping, and Multi-view Re-ranking, to examine their respective contributions. For this analysis, we adopt the GPT-4o backbone and SimCSE retriever, which achieved the strongest overall performance in the main experiments (Table 1). This setting allows us to isolate the effect of each stage under the most competitive configuration.
w/o Stage-1 (Tool Bundle Acquisition). Removing the initial bundle prevents PTR from leveraging historical co-usage patterns, forcing the model to start from unresolved queries without prior tool context. This leads to a clear decline across all datasets, particularly in TRACC (e.g., $0.336\rightarrow 0.235$ on ToolLens), confirming that historical bundle information provides a strong inductive bias for identifying effective tool combinations.
w/o Stage-2 (Functional Coverage Mapping). Without this stage, PTR cannot explicitly align the retrieved tools with query-specific functionalities, resulting in redundant or missing tools. Although the model still benefits from bundle retrieval, TRACC drops considerably (e.g., $0.591\rightarrow 0.552$ on RecTools), showing that this stage is critical for achieving a minimal sufficient tool set by pruning irrelevant tools and surfacing unresolved sub-problems.
w/o Stage-3 (Multi-view Based Re-ranking). Eliminating the re-ranking phase removes the complementary views (semantic, historical, and contextual), causing the model to rely solely on the initial coverage output. The performance gap between this variant and the full model (e.g., $0.566\rightarrow 0.591$ on RecTools) demonstrates that aggregating multiple similarity perspectives refines the final tool set and improves precision.
@@ 5 Related Work
Early tool retrieval methods relied on term-based similarity measures such as BM25 (robertson2009probabilistic) and TF-IDF (sparck1972statistical), which matched queries and tool descriptions through lexical overlap. With the advent of dense retrievers (karpukhin2020dense; guu2020retrieval; xiong2020approximate), neural encoders began to capture semantic relationships between queries and tool documents more effectively. Recent approaches further improve retriever training: Confucius (gao2024confucius) introduces an easy-to-difficult curriculum to enhance LLMs’ understanding of tool usage, while iterative feedback methods refine selection based on tool execution outcomes (wang2023mint; xu2024enhancing). ToolkenGPT (hao2024toolkengpt) represents each tool as a learnable token (“toolken”), enabling tool invocation as a token-generation process. Additionally, research on diversity-aware retrieval (carbonell1998use; gao2024vrsd) improves the coverage and complementarity of retrieved tools. Our work differs in its focus on pre-execution tool-set recommendation: selecting a minimal sufficient subset of tools for a query before any tool is invoked. This regime emphasizes precision prior to execution and is evaluated using both TRACC (which combines sufficiency and minimality) and rank-based metrics. A distinct line of studies addresses execution-time tool use and refinement (schick2023toolformer; cai2024large), including frameworks such as Re-Invoke (chen2024re), DRAFT (qu2024exploration), and TECTON (alazraki2025meta), which optimize invocation success via feedback from real tool executions or trajectories. However, these advantages inherently depend on runtime feedback; converting them into an offline, pre-execution variant removes their core learning signal and yields non-diagnostic surrogates. In this view, PTR can serve as a front-end recommender that supplies a lean, capability-complete tool set to downstream execution planners, while the two paradigms target distinct optimization goals.